OpenStudy (fanduekisses):

If the position of a particle on the x-axis at time t is s(t)=-5^2, then the average velocity of the particle for 0<=t<=3 is?

OpenStudy (misty1212):

\[\frac{f(3)-f(0)}{3-0}\] i think

OpenStudy (misty1212):

i bet it is not really \(s(t)=5^2\) though...

OpenStudy (solomonzelman):

velocity is \(\large\color{slate}{ s'(t) }\)

OpenStudy (solomonzelman):

wait, your s(t) is likely with a typo

OpenStudy (fanduekisses):

oops s(t)=-5t^2

OpenStudy (solomonzelman):

yes, so s'(t)=?

OpenStudy (misty1212):

i am pretty sure it is asking for average velocity on the interval not instantaneous velocity i could be wrong i suppose

OpenStudy (solomonzelman):

\(\large\color{slate}{ s (t) =-5t^2 }\) \(\large\color{slate}{ s '(t) =? }\)

OpenStudy (fanduekisses):

-10t

OpenStudy (solomonzelman):

yes, \(\large\color{slate}{ s'(t) =-10t }\)

OpenStudy (misty1212):

try \[\frac{-5\times 3^2-(-5\times 0^2)}{3-0}\]

OpenStudy (solomonzelman):

the derivative s'(t) is velocity, not s(t). s9t) is the position

OpenStudy (misty1212):

then the AVERAGE velocity of the particle

OpenStudy (solomonzelman):

it would be the average derivative from 0 to 3

OpenStudy (misty1212):

the average derivative?

OpenStudy (solomonzelman):

derivative is the velocity, and you are finding the average velocity.... don't you think so?

OpenStudy (solomonzelman):

I have to leave now... apologize guys

OpenStudy (misty1212):

if you find the average of the derivative over that interval, it is an integral the integral will be the original function evaluate at the endpoints, divide by the length of the interval, it is identical to \[\frac{s(3)-s(0)}{3-0}\]

OpenStudy (fanduekisses):

-15

OpenStudy (misty1212):

\[\frac{-5\times 3^2}{3}=-15\] looks good to me why not?

OpenStudy (fanduekisses):

:)

OpenStudy (misty1212):

it was \(s(t)=-5t^2\) right?

OpenStudy (misty1212):

well in fact solomon was right just unnecessary if you take the average value of the derivative, you get the same thing

OpenStudy (fanduekisses):

yes :)

Parth (parthkohli):

Yup. I was about to argue that there's no need to involve calculus in this because the definition of average velocity is \(\Delta s/\Delta t\) -- nothing instantaneous in that... which is also what you said, so...